Mechanics

A small spherical ball of density p is gently released in a liquid of density d(p>d). Find the initial acceleration of the ball. Post ASAP.

Abc
11 Years agoGrade

2 Answers

Shobhit Varshney
11 Years ago
Hi,

Initially,

the forces acting on the body would be gravity and buoyancy in the opposite directions. find the net force on the body and find the acceleration.

thanks..............................................................................................................................................
Akshat garg
9 Years ago
The forces acting will be buoyant force and weight so W-F=MA. NowF=boyant force=dVg Where v=volumeAlso,W=mgNow,-dvg+mg=ma......(1)We know that mass is density*volume so M=PV Putting in(1)Pvg-dvg=ma(P-d)VG=maVolume is mass/density(m/p)(p-d)mg/p=ma So a is (p-d)g/p=aThank u